# Flavor Equilibration of Supernova Neutrinos: Exploring the Dynamics of Slow Modes

### v1

@@ VI DISCUSSION AND OUTLOOK
We have investigated the phenomenon of flavor equilibration in the context of SFCs within a periodic box for a two-flavor monochromatic neutrino gas. Our findings indicate that a generic equilibration takes place when the densities are averaged over the periodic domain and the angular distributions in the neutrino gas for $0.7\lesssim\alpha\lesssim 1.3$. This leads to a state where the subdominant neutrino species systematically approach near-complete equipartition with slight flavor overconversion than perfect equipartition. The dominant species adjust accordingly to ensure the conservation laws are satisfied. Although near equipartition occurs in both the IO and NO cases, it is generally observed that the results in IO are closer to equipartition than in NO. We have also suggested an empirical formula that can well approximate the coarse-grained flavor conversion probabilities for both $\nu_{e}$ and $\bar{\nu}_{e}$, utilizing the approximate (E-X)LN conservation when the effective vacuum mixing angle is small.
@@ Appendix A Conserved quantities
[…]
@@
$\vec{H}_{\omega}=\pm\omega\begin{pmatrix}\sin{2\theta_{V}}\\ 0\\ \cos{2\theta_{V}}\end{pmatrix}\ ,$ (26)
with $+\omega$ for neutrinos and $-\omega$ for antineutrinos. We assume a two-energy mode system e.g. two possibilities of vacuum frequencies, although our findings carry out straightforwardly to multi-energy calculations. Here, we keep $\theta_{V}$ arbitrarily large not to restrict ourselves to the matter suppressed assumption just yet.
@@
$\vec{P}_{\omega,v_{z},z}^{t_{0}}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}\ ,$ (27)
@@
[…]
@@
[…]
@@
$\partial_{t}\vec{S}_{\bar{\nu}}+\partial_{z}\vec{F}_{\bar{\nu}}=$displaystyle-- $\vec{H}_{\omega}\times\vec{S}_{\bar{\nu}}$ (37)
@@
During the evolution of the system, the parallel component $\langle S_{||}\rangle\equiv\langle\vec{S}\cdot\vec{H}_{\omega}\rangle$ remains a constant of motion. In the special case where the vacuum mixing angle is matter-suppressed $\theta_{V}\ll 1$, the direction of $\vec{H}_{\omega}$ almost aligns with the flavor direction $\hat{z}$. This imposes an approximate conservation of (E-X)LN that relates the equilibrium value of $\langle P_{ee}\rangle$ to that of $\langle\bar{P}_{ee}\rangle$, as described by Eq. 23 in the main text. Throughout this work, the component $\langle S_{||}\rangle$ is conserved to good approximation.

### v2

@@ VI DISCUSSION AND OUTLOOK
We have investigated the phenomenon of flavor equilibration in the context of SFCs within a periodic box for a two-flavor monochromatic neutrino gas. Our findings indicate that a generic equilibration takes place when the densities are averaged over the periodic domain and the angular distributions in the neutrino gas for $0.7\lesssim\alpha\lesssim 1.3$. This leads to a state where the subdominant neutrino species systematically approach near-complete equipartition with slight flavor overconversion than perfect equipartition. The dominant species adjust accordingly to ensure the conservation laws are satisfied. Although near equipartition occurs in both the IO and NO cases, it is generally observed that the results in IO are closer to equipartition than in NO. We have also suggested an empirical formula that can well approximate the coarse-grained flavor conversion probabilities for both $\nu_{e}$ and $\bar{\nu}_{e}$, utilizing the approximate (E-X)LN conservation when the effective vacuum mixing angle is small. Unlike in the case of FFC, there is no clear separation of scales that would allow the disentanglement of advection, collisions, and flavor conversion. Nevertheless, the asymptotic state examined in this study should be interpreted as a quasi-equilibrium state of a local system that approximately conserves lepton number. Our prediction of this quasi-steady state may be useful for schemes that employ relaxation methods to incorporate FC parametrically into hydrodynamical simulations [84, 96, 82, 83].
@@ Appendix A Conserved quantities
$(\partial_{t}+v_{z}\partial_{z})\vec{P}_{\omega,v_{z},z}$ $=$ $\vec{H}_{\omega}\times\vec{P}_{\omega,v_{z},z}+\Big{[}\int d\omega^{\prime}dv_{z}^{\prime}g(\omega^{\prime},v_{z}^{\prime})\vec{P}_{\omega^{\prime},v_{z}^{\prime},z}(1-v_{z}v_{z}^{\prime})\Big{]}\times\vec{P}_{\omega,v_{z},z}\ .$ (25)
@@
$\vec{H}_{\omega}=\pm\omega\begin{pmatrix}\sin{2\theta_{V}}\\ 0\\ \cos{2\theta_{V}}\end{pmatrix}\$, (26)
with $\omega$ for neutrinos and $\omega$ for antineutrinos. We assume a two-energy mode system e.g. two possibilities of vacuum frequencies, although our findings carry out straightforwardly to multi-energy calculations. Here, we keep $\theta_{V}$ arbitrarily large not to restrict ourselves to the matter suppressed assumption just yet.
@@
$\vec{P}_{\omega,v_{z},z}^{t_{0}}=\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}\$, (27)
@@
$\vec{S}_{\nu}$ $\equiv$ $\int dv_{z}g(\omega,v_{z})\vec{P}_{\omega,v_{z},z}h(\omega)\ ,$ (28)
$\vec{S}_{\bar{\nu}}$ $\equiv$ $\int dv_{z}g(\omega,v_{z})\vec{P}_{\omega,v_{z},z}h(-\omega)\ ,$ (29)
@@
$\vec{F}_{\nu}$ $\equiv$ $\int dv_{z}g(\omega,v_{z})v\vec{P}_{\omega,v_{z},z}h(\omega)\ ,$ (32)
$\vec{F}_{\bar{\nu}}$ $\equiv$ $\int dv_{z}g(\omega,v_{z})v\vec{P}_{\omega,v_{z},z}h(-\omega)\ .$ (33)
@@
$\partial_{t}\vec{S}_{\bar{\nu}}+\partial_{z}\vec{F}_{\bar{\nu}}=$ - $\vec{H}_{\omega}\times\vec{S}_{\bar{\nu}}$ (37)
@@
During the evolution of the system, the parallel component $\langle S_{||}\rangle\equiv\langle\vec{S}\cdot\vec{H}_{\omega}\rangle$ remains a constant of motion. In the special case where the vacuum mixing angle is matter-suppressed $\theta_{V}\ll 1$, the direction of $\vec{H}_{\omega}$ almost aligns with the flavor direction $\hat{z}$. This imposes an approximate conservation of (E-X)LN that relates the equilibrium value of $\langle P_{ee}\rangle$ to that of $\langle\bar{P}_{ee}\rangle$, as described by Eq. 23 in the main text. Throughout this work, the component $\langle S_{||}\rangle$ is conserved to a good approximation.